8. A bag contains 3 red balls and 5 black balls. A ball is drawn at random from the bag. What is the probability that the ball drawn is

(i) red?

(ii) not red?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: The total number of balls = No. of red balls + No. of black balls So, the total no. of balls = 5+3 = 8 We know that the probability of an event is the ratio between the no. of favourable outcomes and the total number of outcomes. P(E) = (Number of favourable outcomes/ Total number of outcomes) (i) Probability of drawing red balls = P (red balls) = (no. of red balls/total no. of balls) = 3/8 (ii) Probability of drawing black balls = P (black balls) = (no. of black balls/total no. of balls) = 5/8